A decapeptide (Mol. Wt. 796) on complete hydrolysis gives glycine (Mol. Wt. 75), alanine and phenylalanine. Glycine contributes 47.0% to the total weight of the hydrolysed products. The number of glycine units present in the decapeptide is

Grade:11

1 Answers

Bristi Mukherjee
6 years ago

Decapeptide molecular mass = 796
A decapeptide on hydrolysis give glycine, alanine, phenylalanine.
That means it need 9 water molecule for its hydrolysis.
So the mass of the product = 796 + 9 x 18 = 958
In the total mass 47% = glycine
So the glycine contribution = 47/100 x 958 = 450
Glycine molecular mass = 75
No. of units of glycine = 450/75 = 6